OpenStudy (anonymous):

HELP PLEASE: h(x) = (sin7^(x)) (cos^9 (x)) is an odd or even function

OpenStudy (turingtest):

h(x) = (sin7^(x)) (cos^9 (x)) I assume you meant h(x) = (sin^7(x)) (cos^9 (x)) h(-x)=sin^7(-x)(cos^9(-x) sin(-x)=-sinx and cos(-x)=cosx since sin is raise to an odd power is stays negative so h(-x)=-sin^7(x)cos^9(x)=-h(x) so the function is odd